Four golf balls are to be placed in six different containers. One ball is red; one, green; one, blue; and one, yellow.
a) In how many ways can the four golf balls be placed into six different containers? Assume that any container can contain any number of golf balls (as long as there are a total of four golf balls).
b) In how many ways can the golf balls be placed if container one remains empty?
c) In how many ways can the golf balls be placed if no two golf balls can go into the same container?
d) What is the probability that no two golf balls are in the same container, assuming that the balls are randomly tossed into the containers?
thanks

Answer by edjones(8007)

a)
4 in DISABLED_event_one= 6
3 in 1 and 1 in another: 4C3 * 6P2 = 120
2 in 2: 4C2 * 6C2 = 90
2 in 1 and 1 in 2 others: 4P2 * 6C3 = 240
1 in 4 separate ones: 6P4 = 360
.
6+120+90+240+360=816 ways
.
c)
6P4=360
.
d)
1 * 5/6 * 4/6 * 3/6 = 60/216 = 5/18
.
b) was not clear to me.
.
Please let me know if a) is correct.
.
Ed